OpenStudy (anonymous):

Find two numbers whose difference is 5 and whose product is a minimum

Parth (parthkohli):

What are the 'numbers'? Integers? Natural numbers? Whole numbers?

OpenStudy (anonymous):

(0,5) , (2.5, -2.5) , (7.5, 2.5) , (10, 5)

Parth (parthkohli):

Anyway, there's a logic behind this question. Two negatives multiplied to each other will make a positive. Two positives multiplied to each other will make a positive. A negative and a positive will make a negative.

Parth (parthkohli):

Negative is always less than a positive, so one number must be negative. See what I did there?

Parth (parthkohli):

But as you are given the choices, you may find out their products. \(5 \times 0 = 0\) \(2.5 \times -2.5 = -6.25\) \(7.5 \times 2.5 = 18.75\) \(10 \times 5 = 50\)

Parth (parthkohli):

Which number, according to you, is the least?

OpenStudy (anonymous):

so it has to be (2.5, -2.5) because one has to be negative?

Parth (parthkohli):

Yes, exactly!

OpenStudy (anonymous):

ohhhh! mmmm, i get it :)